OpenStudy (anonymous):

I looking for direction on this question: 1A-6: Express in the form Asin(x+c). For a) sinx +√3cosx & b) sinx-cosx. I'm thinking I need the Sine sum formula: sin(a+b) = sina cosb + sinb cosa, but I haven't trouble getting there.

OpenStudy (anonymous):

To arrive at \[Csin(x+\delta)\] from \[Asinx+Bcosx\] or for this problem \[sinx+\sqrt{3}cosx\] First Obtain A B and C so in this case its \[A = 1;B=\sqrt{3}\] \[C = \sqrt{A^2+B^2}\] \[C = \sqrt{1^2+\sqrt{3}^2} = \sqrt{1+3} = 2; \] Now to find \[\delta\] \[sin(\delta)= B/C\] \[ cos(\delta)= \frac{A}{C}\] In this case \[sin(\delta)=\frac{\sqrt{3}}{2}; \delta =\frac{\pi}{3}\] \[cos(\delta)=\frac{1}{2}; \delta = \frac{\pi}{3}\] or \[tan(\delta) = \sqrt{3}\] So \[\delta = \frac{\pi}{3}\] Now just assemble the equation and \[Csin(x+\delta)\] becomes \[2sin(\frac{\pi}{3}+x)\] This is the basic identity, if you would like a little more rigor there are numerous examples on the web or in textbooks, however this will suffice for rewriting this type of identity.

OpenStudy (anonymous):

Thanks...I was making the problem harder then it had to be.